OpenStudy (anonymous):

a bridge is built in the shape of a paraboolic arc. the bridge arch has a span of 152 feet and a max hight of 25 feet. find the hight of the arc at 20 feet from its center.

OpenStudy (mathmate):

The parabola has a vertex at (0,25), so the equation of the parabola is y = -ax^2+25. Since we know that the arch passes through (76,0), we put 0=-a(76)^2+25 => a=25/76^2 The final equation is then y=-(25/76^2)x^2+25 at x=20, y=-(25/76^2)(20^2)+25=23.27'

OpenStudy (anonymous):

thank you

OpenStudy (mathmate):

yw! :)